You want to dilute and plate an E. coli culture that has 3 *10 8 cells/ml. If you needed to dilute the above culture, starting from the undiluted stock, to a concentration of 500 cells/ml, how could you perform a serial dilution to achieve the desired final concentration?
Answer #1

We have to use the colony forminv unit/ml formula in this situation:

CFU/ml = (No. Of colonies× dilution factor)/ volume plated.

CFU/ml=3 *10^8 cells/ml

No. Of colonies= 500

Volume plated = 1 ml (deally)

3×10^8= (500× dilution factor)/ 1

Dilution factor= 3×10^8/500= 600,000

Dilution = 1/600,000

Dilution 1 from sample: Take 1 ml from sample and dissolved in to 9ml of distilled water=1/(9+1)= 1/ 10

Dilution 2: take 1 ml from dilution 1, and fissolve in 99 ml of water = 1/(99+1)=1/100

Dilution 3: take 1 ml frok dilution 2 and dissolve in 9) ml of water : 1/(99+1)=1/100

Dilution 4: take 1 ml from dilution 3, and dissolve in 5 ml of distilled water=1/(5+1)=1/6

Serial dilution= Dilution (1×2×3×4)= (1/10)×(1/100)×(1/100)×(1/6)= 1/600,000

At last you have to take 1 ml from dilution 4 and then plated on culture plate.